# NEAT: Neighborhood-Guided, Efficient, Autoregressive Set Transformer for 3D Molecular Generation

### v1

@@ Acknowledgements
The financial support received for the Christian Doppler Laboratory for Molecular Informatics in the Biosciences by the Austrian Federal Ministry of Labour and Economy, the National Foundation for Research, Technology and Development, the Christian Doppler Research Association, Boehringer-Ingelheim RCV GmbH & Co KG and BASF SE is gratefully acknowledged.
@@ A.1 Extended Related Works
Diffusion models represent the current state of the art in 3D molecular generation. These models (Ho2020DDPM; Song2021ScoreMatching) operate by progressively corrupting the training data with random noise through a forward diffusion process, and training a neural network to reverse this corruption via step-by-step denoising. During inference, new samples are generated by iteratively applying the learned reverse diffusion process to random noise. Prominent diffusion-based molecular generation models include EDM (Hoogeboom2022EDM), MiDi (Vignac2023MiDi), GeoLDM (Xu2023GeoLDM), GeoBFN (Song2024GeoBFN), and SymDiff (Zhang2025SymDiff). They have been successfully employed in multiple applications, but they do have notable limitations. First, they require the number of atoms in a generated molecule to be pre-specified, which can be addressed using dummy atom tokens. However, this imposes an upper bound on molecule size. Second, diffusion models use computationally expensive message-passing operations at each step of the stochastic differential equation (SDE) integration to capture atom interdependencies. This results in high computational costs during inference.
@@
Flow matching (Lipman2023FlowMatching; Albergo2023NormalizingFlows) offers a conceptually simple alternative to diffusion models and often achieves faster inference due to the need for fewer integration steps during the inference process. It is a simulation-free training framework for generative modeling, where a time-dependent velocity field $\psi_{t}$ is learned to couple a tractable source distribution $p_{0}$ with a target distribution $p_{1}$. Typically, $p_{0}$ is chosen as a standard normal distribution, $p_{0}=\mathcal{N}(\mathbf{0},\mathbf{I})$, and the transport is defined over a time interval $t\in[0,1]$. The transformation of $p_{0}$ into intermediate distributions $p_{t}$ is achieved via the push-forward operation #\#, resulting in a trajectory of probability distributions $p_{t}(\mathbf{x}_{t})$ at time $t$. The flow $\psi_{t}$ is modeled using a parameterized function $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$, which is trained to approximate the velocity field. Sampling from the target distribution $p_{1}$ is performed by drawing samples $\mathbf{x}_{0}\sim p_{0}$ and integrating the ordinary differential equation (ODE) $d\mathbf{x}_{t}=\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$.
@@
Lipman2023FlowMatching introduced simulation-free training by constructing conditional probability paths. A widely used choice for these paths is the linear interpolant (Albergo2023NormalizingFlows), defined as $\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}$, with the conditional target velocity $\mathbf{v}_{t}=\mathbf{x}_{1}-\mathbf{x}_{0}$. The training objective, known as conditional flow matching (CFM), minimizes the $l_{2}$ regression loss: $\mathcal{L}_{\text{CFM}}=\mathbb{E}[\lVert\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-\mathbf{v}_{t}\rVert]$. Flow-matching-based molecular generation models include EquiFM (Song2023EquiFM) and FlowMol3 (Dunn2025FlowMol). Independent coupling of $p_{0}$ and $p_{1}$ can result in high transport costs and intersecting conditional paths. These issues can be alleviated by employing optimal transport couplings between the source and target distributions, thereby improving convergence and enhancing flow performance (tong2024improving). In molecular generation, this has been approached, e. g., through rigid alignment of noisy and true coordinates, combined with solving the assignment problem to minimize transport costs (Dunn2025FlowMol).
A.2 Implementation Details
@@
Training was performed on a Rocky Linux 9.4 system with an NVIDIA RTX Pro 6000 Blackwell Graphics Processing Unit with 96 GB GDDR7 VRAM and an AMD EPYC 9555 64-Core Processor. We use CUDA 13.0 for GPU acceleration. Our model is implemented in Python (v3.11.13) with PyTorch (v2.8.0), PyTorch Geometric (v2.7.0) (fey2019pyg), and Pytorch Lightning (v2.5.5) (Falcon_PyTorch_Lightning_2025). Model training was monitored with Tensorboard (v2.20.0). For chemical data processing we employed RDKit (v2025.9.1) (rdkit_2025). For visualization, we used Py3Dmol (v2.5.3), plotting was performed with matplotlib (v3.10.7).
@@
The basis of our implementation is the NanoGPT implementation of Karpathy2025nanoGPT, which provides a concise implementation of the GPT-2 (radford2019language) model. We do not use bias in the layers of the transformer module. For positional encoding of the Cartesian coordinates, we employ Fourier features (tancik2020fourier) as implemented in (wang2025simplefold). We experimented with qk-normalization and rotary position embeddings(su2024roformer), but did not find improved results. We found additive pooling the most promising pooling strategy, but also tried mean and attention pooling. We also tried local attention, defined by a radius graph, instead of full attention over all atom nodes, which led to performance degradation. Inclusion of charged states into the atom type vocabulary did not lead to model improvement. Since our model architecture is not $SE(3)$-equivariant, we augment data samples by random rotation, which stabilizes the training. We experimented with random translations as well, however, we did not find an increase in model performance. The flow matching head is the same adaptive layer-norm MLP as used in Li2024AutoregressiveImageGenerationWithoutVQ, which was also used in the QUETZAL model (Cheng2025Quetzal) for implementation of the diffusion head.
@@
The splitting procedure in Algorithm 1 is governed by the hyperparameters $\beta$ and $\gamma$ (see Section 3). Figure 4 reports the absolute and relative (w.r.t. the original graphs) source and target set sizes for a batch of 25,000 training graphs. Across all runs, we used $\beta=1.5$ and $\gamma=0.55$. These settings produced an approximately uniform source set distribution with a slight bias toward smaller sets. The resulting target sets contain on average $\sim 18\%$ of the nodes of the original graphs. We also evaluated source-set distributions emphasizing mid-sized to larger sets, but near-uniform sampling performed best.
Figure 4: Relative and absolute size distribution of the source and target set w.r.t. the original graphs. Sampling was performed with $\beta=1.5$ and $\gamma=0.55$.
@@
For training the flow matching head, we sample random time steps analog to the SimpleFold model (wang2025simplefold) from the following log-normal distribution:
@@
where we set $m=0.8$ and $s=1.7$. With these settings, time steps are sampled more densely around $t=1$. We found that this sampling schema slightly improved training above time step sampling from a uniform distribution. For each conditional path, we sampled 16 random time steps. While we found resampling to be benefitial, using more than 16 time steps per path did not have an impact on model convergence. We found Euler integration with equidistant time steps to be sufficient for integration of the flow ODE. We also tried a Runge-Kutta-4 and an adaptive time step Dormand-Prince integrator, but could not find improved performance. Using endpoint parametrization for construction of the conditional paths, like done in Dunn2025FlowMol, led to performance degradation.
@@
The hyperparameters of our best performing model are summarized in Table 2. The loss function was minimized with the AdamW (loshchilov2019decoupled) optimizer using a learning rate of $1\cdot 10^{-4}$, parameters $\beta_{1}=0.9$ and $\beta_{2}=0.95$, and a weight decay of $1\cdot 10^{-6}$. We further used gradient clipping to $1.0$ of the gradient norm. We experimented with different learning rates and found that increasing the learning rate led to exploding gradients. We implemented a cosine annealing learning rate schedule with 50 linear warm-up epochs and set the minimum learning rate to 10% of the initial learning rate after $10,000$ epochs. Model training took $96$ hours, given the above hardware specifications.
Table 2: Hyperparameters of the best performing model on the QM9 dataset
@@
learning rate $10^{-4}$
learning rate decay number of epochs 10,000
learning rate minimum $10\%$
learning rate warm-up number of epochs 50
time step resampling 16
@@
A.3 Additional Results & Comparisons

### v2

@@ Acknowledgements
The authors thank Leo Gaskin for testing the code and Steffen Hirte for proofreading the manuscript. The financial support received for the Christian Doppler Laboratory for Molecular Informatics in the Biosciences by the Austrian Federal Ministry of Labour and Economy, the National Foundation for Research, Technology and Development, the Christian Doppler Research Association, Boehringer-Ingelheim RCV GmbH & Co KG and BASF SE is gratefully acknowledged.
@@ A.1 Extended related works
Diffusion models (Ho et al., 2020; Song et al., 2021) operate by progressively corrupting the training data with random noise through a forward diffusion process, and training a neural network to reverse this corruption via step-by-step denoising. During inference, new samples are generated by iteratively applying the learned reverse diffusion process to random noise. Prominent diffusion-based molecular generation models include EDM (Hoogeboom et al., 2022), MiDi (Vignac et al., 2023), GeoLDM (Xu et al., 2023), GeoBFN (Song et al., 2024), and SymDiff (Zhang et al., 2025). They have been successfully employed in multiple applications, but they do have notable limitations. First, they require the number of atoms in a generated molecule to be pre-specified, which can be addressed using dummy atom tokens. However, this imposes an upper bound on molecule size. Second, diffusion models use computationally expensive message-passing operations at each step of the stochastic differential equation (SDE) integration to capture atom interdependencies. This results in high computational costs during inference.
@@
Flow matching (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023) offers a conceptually simple alternative to diffusion models and often achieves faster inference due to the need for fewer integration steps during the inference process. It is a simulation-free training framework for generative modeling, where a time-dependent velocity field $\psi_{t}$ is learned to couple a tractable source distribution $p_{0}$ with a target distribution $p_{1}$. Typically, $p_{0}$ is chosen as a standard normal distribution, $p_{0}=\mathcal{N}(0,I)$, and the transport is defined over a time interval $t\in[0,1]$. The transformation of $p_{0}$ into intermediate distributions $p_{t}$ is achieved via the push-forward operation #\#, resulting in a trajectory of probability distributions $p_{t}(\mathbf{x}_{t})$ at time $t$. The flow $\psi_{t}$ is modeled using a parameterized function $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$, which is trained to approximate the velocity field. Sampling from the target distribution $p_{1}$ is performed by drawing samples $\mathbf{x}_{0}\sim p_{0}$ and integrating the ordinary differential equation (ODE) $\mathrm{d}\mathbf{x}_{t}=\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$.
@@
Lipman et al. (2023) introduced simulation-free training by constructing conditional probability paths. A widely used choice for these paths is the linear interpolant (Albergo and Vanden-Eijnden, 2023), defined as $\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}$, with the conditional target velocity $\mathbf{v}_{t}=\mathbf{x}_{1}-\mathbf{x}_{0}$. The training objective, known as conditional flow matching (CFM), minimizes the $l_{2}$ regression loss: $\mathcal{L}_{\text{CFM}}=\mathbb{E}[\lVert\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-\mathbf{v}_{t}\rVert]$. Flow matching-based molecular generation models include EquiFM (Song et al., 2023) and FlowMol3 (Dunn and Koes, 2025). Independent coupling of $p_{0}$ and $p_{1}$ can result in high transport costs and intersecting conditional paths. These issues can be alleviated by employing optimal transport couplings between the source and target distributions, thereby improving convergence and enhancing flow performance (Tong et al., 2024). In molecular generation, this has been approached, e. g., through rigid alignment of noisy and true coordinates, combined with solving the assignment problem to minimize transport costs (Dunn and Koes, 2025).
A.2 Data processing details
The QM9 dataset was retrieved from the figshare repository of Ramakrishnan et al. (2019) and initially comprised 133,885 molecules. The repository additionally provides a list of 3,054 molecule IDs for which OpenBabel failed to infer correct bonds from the XYZ geometry files. Following prior work (Hoogeboom et al., 2022; Cheng et al., 2025), we excluded these molecules. Because the dataset is distributed as XYZ files without explicit bond annotations, and our neighborhood-guidance strategy requires bond information, we reconstructed bonds for each molecule using RDKit (MolFromXYZBlock together with rdDetermineBonds). Bond assignment failed for 6810 molecules, yielding a processed dataset of 124,021 small molecules. We randomly split these into 100,000/11,619/12,402 molecules for the training/validation/test sets, respectively. While some prior works on 3D molecular generation obtained QM9 via the MoleculeNet project (Wu et al., 2018) rather than the original repository (Ramakrishnan et al., 2019), we decided against using the MoleculeNet files. Although they include bond information, the procedure used to determine bonds is not documented, and a non-negligible fraction of molecules in that distribution are charged. This conflicts with the original QM9 specification (Ramakrishnan et al., 2014), which describes the dataset as comprising neutral molecules with up to nine heavy atoms (C, N, O, F), excluding hydrogen atoms.
The GEOM-Drugs dataset was obtained from the FlowMol project’s data repository (Dunn and Koes, 2025) as distributed by Koes (2024). The files provide a predefined train/validation/test split totaling 304,313 molecules, which we adopted unchanged. We removed 35 molecules that failed RDKit sanitization. For molecules comprising multiple disconnected fragments, we retained only the largest fragment by atom count. Following Vignac et al. (2023), we used up to five low-energy conformations per molecule. After preprocessing, the dataset contained 243,541/30,430/30,432 molecules and 1,172,436/146,394/146,788 conformations in the train/validation/test splits, respectively.
A.3 Implementation details
@@
Training was performed on a Rocky Linux 9.4 system with an NVIDIA RTX Pro 6000 Blackwell GPU with 96 GB GDDR7 VRAM and an AMD EPYC 9555 64-Core Processor. We use CUDA 13.0 for GPU acceleration. Our model is implemented in Python (v3.11.13) with PyTorch (v2.9.0), PyTorch Geometric (v2.7.0) (Fey and Lenssen, 2019), and Pytorch Lightning (v2.5.5) (Falcon and the PyTorch Lightning team, 2025). Model training was monitored with Tensorboard (v2.20.0). For chemical data processing we employed RDKit (v2025.9.1) (Landrum, 2025). For visualization, we used Py3Dmol (v2.5.3) and matplotlib (v3.10.7).
@@
The basis of our implementation is the NanoGPT implementation of Karpathy (2025), which provides a concise implementation of the GPT-2 (Radford et al., 2019) model. We do not use bias in the normalization layers and feed-forward MLPs of the transformer module. For positional encoding of the Cartesian coordinates, we employ Fourier features (Tancik et al., 2020) as implemented in Wang et al. (2025a). We experimented with qk-normalization and rotary position embeddings (Su et al., 2024), but did not find improved results. We found additive pooling to be the most effective pooling strategy, but also tried mean and attention pooling. We also tried local attention, defined by a radius graph, instead of full attention over all atom nodes, which led to performance degradation. Inclusion of charged states into the atom type vocabulary did not lead to model improvement. Since our model architecture is not $SE(3)$-equivariant, we augment data samples by random rotation, which stabilizes the training. We also experimented with random translations, but did not observe an increase in model performance. The flow matching head is the same adaptive layer-norm MLP as used in Li et al. (2024), which was also used in the QUETZAL model (Cheng et al., 2025) for implementation of the diffusion head.
@@
The splitting procedure in Algorithm 1 is dependent on the hyperparameters $\beta$ and $\gamma$ (see Section 3). Figure 5 and Figure 6 report the absolute and relative (w.r.t. the original graphs) source and target set sizes for a batch of 25,000 randomly sampled training graphs. Across all runs, we used $\beta=1.5$ and $\gamma=0.45$. These settings produced an approximately uniform source set distribution with a slight bias toward smaller sets (corresponding to relative size). The resulting target sets contain on average $\sim 18\%$ of the nodes of the original graphs. We also evaluated source-set distributions emphasizing mid-sized to larger sets, but near-uniform sampling performed best.
Figure 5: Relative and absolute size distribution of the source and target set w.r.t. the original graphs in the QM9 dataset. Sampling was performed with $\beta=1.5$ and $\gamma=0.45$.
Figure 6: Relative and absolute size distribution of the source and target set w.r.t. the original graphs in the GEOM dataset. Sampling was performed with $\beta=1.5$ and $\gamma=0.45$.
@@
For training the flow matching head, we sample random time steps analog to the SimpleFold model (Wang et al., 2025a) from the following log-normal distribution:
@@
where we set $m=0.8$ and $s=1.7$. With these settings, time steps are sampled more densely around $t=1$. We found that this sampling schema slightly improved training above time step sampling from a uniform distribution. For each conditional path, we sampled four random time steps, following Li et al. (2024). Using endpoint parametrization for construction of the conditional paths, like done in Dunn and Koes (2025), led to performance degradation.
Sampling from the flow model
We found Euler integration with $N=60$ equidistant time steps to be best for integration of the flow ODE for the model trained on QM9 data. We also tried a Runge-Kutta-4 and an adaptive time step Dormand-Prince integrator, but could not find improved performance. For the model trained on GEOM, best results could be obtained by using the Euler-Maruyama integrator, which is described in the following. Sampling from a flow-based generative model can be performed by integrating the ODE $\mathrm{d}\mathbf{x}_{t}=\mathbf{v}_{\theta}(\mathbf{x}_{t},t)\,\mathrm{d}t$. Beyond deterministic sampling, the velocity field can be used to simulate the following reverse-time stochastic differential equation (SDE) (Ma et al., 2024), thereby yielding a stochastic generative model:
$\mathrm{d}\mathbf{x}_{t}=\mathbf{v}_{\theta}(\mathbf{x}_{t},t)\mathrm{d}t+\frac{1}{2}w(t)\mathbf{s}_{\theta}(\mathbf{x}_{t},t,c)\mathrm{d}t+\sqrt{\tau\cdot w(t)}\mathrm{d}\mathbf{\bar{W}}_{t}$ (8)
where $\mathbf{\bar{W}}_{t}$ is a reverse-time Wiener process, $w_{t}$ is the diffusion coefficient, and $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)=\nabla\log p_{t}(\mathbf{x}_{t})$ is the Stein score function. We employ the same Euler-Maruyama integrator as described in Wang et al. (2025a). The score $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)=\frac{t\mathbf{v}_{\theta}(\mathbf{x}_{t},t)-\mathbf{x}_{t}}{1-t}$ can be derived from the velocity and the diffusion coefficient is calculated as $w(t)=\frac{2(1-t)}{t+\eta}$, where $\eta=10^{-3}$ for numerical stability. In all of our experiments we use $\tau=0.3$, $N=60$ integration steps, and equidistant time step spacing.
@@
Table 3 and Table 4 summarize the hyperparameters of our best-performing models on QM9 and GEOM. The models have 282M (QM9) and 165M (GEOM) parameters. Results from the hyperparameter screening experiments on GEOM are given in Table 5. The loss function was minimized with the AdamW (Loshchilov and Hutter, 2019) optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We further used gradient clipping to $1.0$ of the gradient norm. We implemented a cosine annealing learning rate schedule with linear warm-up epochs and set the minimum learning rate to 10% of the initial learning rate. Model training took 60 and 106 hours on QM9 and GEOM, respectively, given the hardware specifications listed above.
Table 3: Hyperparameters of the best performing NEAT model trained on the QM9 dataset.
@@
$\gamma$ 0.45
integrator Euler
learning rate $1.0\times 10^{-4}$
learning rate decay number of epochs (cosine annealing scheduling) 10,000
learning rate minimum (cosine annealing scheduling) $10\%$
learning rate warm-up number of epochs (cosine annealing scheduling) 50
time step resampling 4
@@
Table 4: Hyperparameters of the best performing NEAT model trained on the GEOM dataset.
Parameter Value
batch size 512
$\beta$ 1.5
bias true
dropout 0.1
epochs 1,000
flow matching head hidden dimension 1,536
flow matching head layers 6
flow matching noise standard deviation $\sigma$ 2.5
$\gamma$ 0.45
integrator Euler-Maruyama
learning rate $2.5\times 10^{-5}$
learning rate decay number of epochs (cosine annealing scheduling) 1,000
learning rate minimum (cosine annealing scheduling) $10\%$
learning rate warm-up number of epochs (cosine annealing scheduling) 10
time step resampling 4
time step sampling logit normal
transformer heads 12
transformer hidden dimension 768
transformer layers 12
vocabulary size 17
weight decay $10^{-6}$
Table 5: Hyperparameter screening results on the GEOM dataset. The configuration selected for the final model is highlighted in italics, chosen to balance optimal performance metrics with model size.
transformer heads transformer layers transformer dimension FM head dimension model parameters $\times 10^{6}$ atom stable lookup valid lookup unique xyz2mol valid xyz2mol unique
8 8 512 1536 104.32 79.4 93.8 93.6 66.3 66.0
10 10 640 1536 128.62 79.8 97.5 97.3 68.0 67.9
12 12 768 1536 164.72 81.1 97.9 97.4 74.2 73.7
14 14 896 1536 214.97 80.7 97.4 97.3 72.6 72.5
16 16 1024 1536 281.72 80.6 97.0 96.9 71.8 71.8
18 18 1152 1536 367.40 81.6 96.7 96.6 69.6 69.6
12 12 768 512 95.20 79.2 96.4 96.3 68.5 68.4
12 12 768 1024 121.57 80.5 98.2 98.0 70.3 70.1
12 12 768 1536 164.72 81.1 97.9 97.4 74.2 73.7
12 12 768 2048 224.65 82.0 98.0 97.9 74.2 74.1
12 12 768 2560 301.35 81.9 99.0 98.9 77.7 77.6
A.4 Observations on evaluation metrics
Hoogeboom et al. (2022) introduced a pipeline for generating molecular structures from atomic point clouds by using a look-up table of interatomic distances for single, double, and triple bonds. A bond is established if the interatomic distance is smaller than the corresponding table value plus a defined tolerance margin, starting with the shortest bond type and progressing sequentially to larger bond types. If no bond type matches the distance criteria, the atoms remain unbonded. Atom stability within this framework is defined by adherence to the octet or duet rule based on the assigned bonds. Molecular stability is achieved when all constituent atoms are stable, and molecular validity is determined by whether the molecule can be converted into a valid SMILES string after sanitization using RDKit.
A notable limitation of this pipeline is that it processes interatomic distances in a sequential manner, which does not guarantee convergence to an optimal solution. Moreover, this pipeline does not consider aromatic bonds, which have lengths between single and double bonds. Consequently, aromatic molecules are often falsely categorized as unstable. For example, evaluating benzene with this pipeline results in 0% atom stability. This occurs because all bonds in benzene are incorrectly assigned as ”double” bonds, leading to a chemically invalid configuration where each carbon atom has five bonds (two double bonds to neighboring carbon atoms and one single bond to a hydrogen atom). Given the high prevalence of aromatic compounds in the GEOM-Drugs dataset, this issue leads to a surprisingly low molecular stability on the GEOM-Drugs test set and on the sets of molecules generated from models trained on GEOM-Drugs.
Another peculiarity of the pipeline is that the code used in Hoogeboom et al. (2022) and related works (Vignac et al., 2023; Cheng et al., 2025) reduces all assigned bonds to single bonds when evaluating the GEOM-Drugs dataset. This simplification allows many SMILES strings to pass the RDKit sanitization test, even though they contain carbon atoms with incorrect multiplicities. To ensure fair comparison with other models, we have adhered to this widely used pipeline without modifications.
We also evaluate models using the xyz2mol metric, which tests whether generated 3D atomic point clouds can be converted into valid SMILES after bond assignment via RDKit’s rdDetermineBonds. We would like to point out that this metric is sensitive to the RDKit version used. We observe decreases of up to $5\%$ with RDKit versions later than v2023.3.3, likely due to stricter bond assignment. We use RDKit v2025.9.1 throughout this work, which was the latest available update at project start. We recomputed xyz2mol metrics where model weights or pre-generated samples were available. Literature-reported xyz2mol results with undocumented RDKit versions are marked accordingly. While xyz2mol appears more robust than EDM lookup metrics, we advocate for alternative evaluation pipelines to improve the accuracy and reliability of molecular generator assessment.
A.5 Additional results
Molecular properties of generated molecules
We compared the generated molecules to the training set across eight molecular properties: molecular weight, Crippen logP, topological polar surface area (TPSA), ring count, fraction of rotatable bonds, fraction of heteroatoms, fraction of halogen atoms, and fraction of stereocenters (Figure 7). Overall, the distributions of these properties are broadly similar between the two sets. Closer inspection indicates that the generated molecules exhibit slightly broader distributions for molecular weight, logP, and TPSA. The distributions of ring count, fraction of rotatable bonds, and fraction of heteroatoms are slightly skewed toward lower values in the generated set. This skewness is also observed for the fractions of halogen atoms and stereocenters, and is more pronounced for these two properties.
Figure 7: Comparison of generated molecules to the training set across eight molecular properties: molecular weight, Crippen logP, topological polar surface area, ring count, fraction of rotatable bonds, fraction of heteroatoms, fraction of halogen atoms, and fraction of stereocenters. Each panel shows overlaid histograms (training: red; generated: blue), reported as frequency in percent. Fractional properties are normalized by heavy atom count. Integer-valued properties use integer-centered bins.
Examples of generated molecules
Figure 8, 9, 10, and 11 show 2D and 3D plots of randomly selected molecules generated by NEAT trained on the QM9 and on the GEOM-Drugs datasets, respectively.